OpenStudy (anonymous):

Rosie had $30 in her purse after she baby-sat for the Nolan children this afternoon. Since she was running late, it cost her $5 to take a cab so she could arrive at the Nolan house by 2 P.M. If she made $25 baby-sitting but spent $1 on bus fare back home, how much money was in her purse before she baby-sat? A. $35 B. $26 C. $20 D. $11

OpenStudy (anonymous):

D $11 11 - 5 = $6 6 + 25 = $31 31 - 1 = $30

OpenStudy (anonymous):

You've already chosen the best answer. 0Hopes this helps http://www.wolframalpha.com/input/?i=Multiply+%28x%2B2%29%282x%5E2-x%2B4%29+

OpenStudy (anonymous):

hahaha jp

OpenStudy (anonymous):

I havent chosen the best answer yet

OpenStudy (anonymous):

choose mine lol

OpenStudy (anonymous):

U havent answered the question that is wat i sended u did u understand it?

OpenStudy (anonymous):

probrably d=11$

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Chlorophyll wat do u think it is

OpenStudy (anonymous):

whose Chlorophyll

OpenStudy (anonymous):

he is @Chlorophyll

OpenStudy (anonymous):

x - 6 + 25 = 30 => x = 30 + 6 -25 = $11

OpenStudy (anonymous):

Does it make sense to you?

OpenStudy (anonymous):

yes chlorophyll thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Thanks Chlorophyll but, To be fair im giving it to @tfitz17

OpenStudy (anonymous):

caz she answered first i just needed another opinion.

OpenStudy (anonymous):

I don't mind at all, as long as the asker understand :)

OpenStudy (anonymous):

thats the spirit @chlorophyll